What is the solution to 2x=8 and x=6+y using substitution
Answer by Alan3354(69443)

What is the solution to 2x=8 and x=6+y using substitution
============
2x = 8 --> x = 4
x = 6 + y
Sub for x
4 = 6 + y
y = -2